5. If a firms total cost is TC-36+402 Which output minimizes the firms average cost? А. 3 B. 4 C. 6 D. 12
Answer #1

Answer 5

The correct answer is (A) 3

Average Cost(AC) = TC/Q = (36 + 4Q2)/Q = 36/Q + 4Q

Hence we have to Minimize : AC = 36/Q + 4Q

First order condition :

dAC/dQ = 0 => -36/Q2 + 4 = 0 => 4 = 36/Q2 => Q2 = 9

=> Q = 3

So, Output that minimizes Average Cost is 3 units

Hence, the correct answer is (A) 3